OpenStudy (anonymous):

0= -x^2 + 14x - 49, i've gotten this far but now what?

OpenStudy (lgbasallote):

wait...are you sure it's -49? not positive?

OpenStudy (kinggeorge):

\[0=-x^2+14x-49=-(x^2-14x+49)\]\[0=-((x-7)(x-7))\]\[0=(x-7)(x-7)\]Can you finish solving the problem from there?

OpenStudy (anonymous):

oh ok thanks

OpenStudy (kinggeorge):

you're welcome